In a Math test, the mean score is 45 and the standard deviation is 4. Assuming normality, what is the probability that a score picked at random will lie below 45

1
Expert's answer
2022-02-28T14:17:46-0500

Since "\mu" =45 and "\sigma"=4 we have:

P(X<45) = P(X-"\mu" < 45-45) = P("\frac{x-\mu}\sigma" < "\frac{45-45}{4}")


Since


Z = "\frac{x-\mu}{\sigma}" = "\frac{45-45}{4}" = 0


Then


P(X<45) = P(Z<0)

From the normal distribution table, we find that:

P(Z<0) = 0.5

Therefore, we conclude that the probability that a score picked at random will lie below 45 is 0.5